# Emotional Support Evaluation Framework via Controllable and Diverse Seeker Simulator

### v1

@@ Abstract
Emotional Support Evaluation Framework via Controllable and Diverse Seeker Simulator
Chaewon Heo1  Cheyon Jin1  Yohan Jo1thanks: Corresponding author 1Graduate School of Data Science, Seoul National University {heorshey99, cheyonjin, yohan.jo}@snu.ac.kr
@@ 1 Introduction
Currently, the most prevalent evaluation strategy involves using help-seeker simulators to interact with supporter models, and the resulting dialogues are assessed based on performance metrics such as empathy and fluency Zhao et al. (2024). However, existing seeker simulators compromise evaluation validity due to two critical limitations. First, they fail to reflect the diversity of real-world seekers, representing only a narrow subset of seeker behaviors. Existing simulators primarily portray overly cooperative, "easy" seekers, failing to capture the vast spectrum of real-world behaviors such as advice resistance Yaman (2021) or minimal utterances. Second, existing seeker simulators lack fine-grained controllability. A robust evaluation framework must allow researchers to target specific seeker populations, as this enables the identification of which aspects of a supporter model fail for particular seeker groups, providing actionable signals for model improvement.
To address these shortcomings, we present a highly controllable seeker simulator designed to take specific seeker profiles and simulate their corresponding behaviors. As a foundational component, we define the seeker profile as a combination of nine psychological and linguistic features (e.g., resistance levels and coping strategies). To faithfully reflect a broad array of real-world behaviors, we leverage data from Reddit online support groups as the primary training source for our simulator.
To effectively control seeker profile defined by nine features, we employ a Mixture-of-Experts (MoE) architecture. This architecture dynamically assigns weights to specialized adapters based on the input seeker profile, enabling distinct behavioral patterns associated with different characteristics to be accurately reflected. Validation results highlight the effectiveness of our simulator across three key dimensions: (1) superior profile adherence, successfully maintaining consistent seeker traits throughout the interaction; (2) high behavioral diversity, capturing a wider array of help-seeking patterns compared to existing baselines; and (3) higher fidelity, as confirmed by expert evaluations indicating that the generated responses more closely resemble real-world emotional support interactions.
Using our controllable seeker simulator, we evaluate supporter models through interactive multi-turn dialogues, followed by automatic assessment using emotional support and general conversational skill metrics (see Figure 1). Results indicate that supporter models exhibit varying levels of performance degradation across various seeker populations. Our framework allows researchers to flexibly define seeker populations by selecting appropriate combinations of profile features, enabling targeted and diagnostic evaluation under diverse scenarios.
@@ 2.1 Evaluation of Emotional Support Models
The evaluation of emotional support models has evolved from surface-level fluency to interactional supportive efficacy. Early research focused on evaluating whether generated empathetic responses exhibited emotionally appropriate expressions Santhanam and Shaikh (2019), using lexical metrics (e.g., BLEU, Perplexity) against reference dialogues such as EmpatheticDialogues Rashkin et al. (2019). Since these metrics often fail to capture supportive effectiveness, later work adopted ESConv Liu et al. (2021) as evaluation scenarios and performed human evaluation, where evaluators role-play seekers and rate support quality using predefined criteria (e.g., Identification, Empathy) Zhao et al. (2023); Cheng et al. (2022). However, these human role-play evaluations are constrained by predefined scenarios and fail to fully reflect the dynamics of real-world support interactions, while also being costly to conduct at scale. These limitations have driven recent evaluation toward simulator-based evaluation frameworks, where seeker simulators interact with supporter models and automated evaluation is performed on the generated dialogues.
@@ 2.2 Seeker Simulations in Emotional Support
Prior studies have adopted diverse approaches to implement seeker simulators. Prompt-based simulators are given input including specific characteristics, such as Big Five personality traits, or situational contexts involving various mental health issues Madani and Srihari (2025); Qiu and Lan (2024). Beyond simple prompting, some work employs additional control mechanisms, such as explicit state tracking or memory modules, to directly regulate the seeker’s internal states Yang et al. (2025); Wang et al. (2025). Another work trains seeker simulators on emotional support datasets using profile cards with demographic attributes, enabling the simulator to learn seeker behaviors associated with each profile Zhao et al. (2024).
Despite conditioning on different seeker prompts, existing simulators typically inherit the base language model’s verbose and cooperative interaction bias. Moreover, even learned simulators tend to reflect only surface-level profile attributes, failing to induce deeper psychological traits across seeker types. Our simulator design addresses these limitations.
@@ 3.1 Data Collection and Preprocessing
We curated a dataset of real-world emotional support interactions from Reddit, specifically targeting subreddits with over 500,000 members, such as r/offmychest and r/mentalhealth. Reddit serves as an ideal data source for training a feature-controllable simulator due to its in-the-wild nature. It encompasses an exceptionally broad spectrum of informal, diverse, and raw human expressions that are often absent in curated datasets.
Unlike traditional approaches that rely on heavily filtered datasets, we intentionally retained informal and raw elements. We included a wide variance in utterance lengths and aggressiveness, which directly addresses the limitations of prior simulators that portray only highly articulate and cooperative seekers.
@@ 3.2 Feature Taxonomy for Seeker Profiles
To represent the multifaceted characteristics of seekers’ conversational behavior, we define a comprehensive taxonomy consisting of nine distinct features. These are categorized into psychological and linguistic attributes. Detailed category definitions for each feature are provided in the Appendix B
@@ 3.2.1 Psychological Features
We incorporate several psychological features to model seeker behavior. The main coping strategy categorizes seekers into four types, grounded in the multidimensional coping framework Endler and Parker (1994): problem_focused, emotion_focused, avoidant, and maladaptive_behavior. The utterance style captures stylistic variation in seeker expression using three categories (plain, upset, verbose), adapted from the PATIENT-$\Psi$ framework Wang et al. (2024). To model interactional dynamics, we define a three-level ordinal resistance level based on psychological reactance theory Brehm and Brehm (1981), and a three-level engagement level grounded in prior work on client participation in therapeutic settings Holdsworth et al. (2014). Finally, we include a four-level self-disclosure level derived from Social Penetration Theory Altman and Taylor (1973), as well as seeker reaction proportions following prior work on modeling client reactions in emotional support dialogues Li et al. (2023).
@@ 3.2.2 Linguistic Features
Complementing these psychological features, we define three linguistic attributes that capture diverse structural patterns. The verbosity level is defined as a five-level feature based on the distribution of seeker utterance lengths. We additionally include a binary user profanity flag, detected using the profanity_check library. Finally, the total dialogue turns level serves as a structural feature defining the overall length of the interaction.
@@ 3.3 Feature Annotation and Validation
To apply the feature taxonomy defined in Section 3.2 to our corpus, we employed a hybrid annotation framework that combines LLM-based tagging for psychological features and rule-based extraction for linguistic features.
@@ 3.3.1 Psychological Feature Annotation
To verify the reliability of the LLM-based feature tagging, we conducted a validation study on a sample of 60 dialogues. Inter-annotator agreement (percent agreement) averages 0.57 across psychological features, while human–LLM alignment achieves 0.84 accuracy, indicating reliable alignment. Detailed validation procedures and feature-level validation results are provided in the Appendix D.
@@ 3.3.2 Linguistic Feature Extraction
Linguistic features describe the measurable characteristics of the dialogue and are extracted using rule-based methods. These include the verbosity level derived from seeker utterance length, the profanity flag extracted using the profanity-check library, and the total dialogue turns level derived from the total number of turns. Detailed extraction rules are provided in the Appendix E.
@@ 3.4 Seeker Profile Construction
We construct a final seeker profile by integrating all annotated features defined in the previous sections with the seeker’s main problem. Formally, a seeker profile is represented as:
@@
where each $f_{i}$ denotes an annotated seeker feature constructed in section 3.2 and $m$ represents the seeker’s main problem, both expressed in natural language. An example seeker profile is provided in the Figure 8. The resulting dataset consists of 11,066 unique seeker profiles, split into 8,868 training, 1,094 validation, and 1,104 test profiles. The seeker’s main problem is a summary of the first seeker utterance in the dialogue (i.e., original post) that describes the seeker’s situation. The summaries are generated by GPT-4o-mini OpenAI (2024), with detailed prompts provided in Table 7.
Figure 2: Overview of the MoLA architecture. A frozen SFT backbone is augmented with multiple low-rank expert adapters at each linear layer. A shared routing network produces a dialogue-level routing vector $\boldsymbol{\alpha}$, which controls expert activation consistently across all layers.
@@ 4 Seeker Simulator Training
To simulate the diverse behaviors of seekers in emotional support conversations, we employ a two-stage training framework: (1) Supervised Fine-Tuning (SFT) to establish a general conversational backbone, and (2) Mixture-of-Experts (MoE) training to enable controllable, attribute-specific generation.
@@ 4.1 Supervised Fine-Tuning
The first stage establishes a linguistic foundation by capturing general conversational patterns of seekers. We use Llama-3-8B-Instruct as the base model and apply LoRA (we set rank $r=16$) to all linear layers. The model is trained to predict the next seeker utterance, given a seeker profile and dialogue history as a next-token prediction task. Upon convergence, the LoRA adapters are merged into the base model weights. This merged model is then entirely frozen, serving as a stable backbone for the subsequent expert-based training stage.
@@ 4.2 MoE Training with Behavioral Routing
To address this challenge, we introduce an explicit routing mechanism that decouples feature control from language modeling. Rather than expecting the model to infer seeker features from text, we provide a structured feature vector to control expert selection through a soft MoE formulation. An overview of the proposed framework is illustrated in Figure 2.
@@ 5.1.1 Experimental Setup
  • Zero-shot Base Models: GPT-4.1-mini, Llama-3-8B-Instruct Touvron et al. (2024), and Qwen-2.5-14B-Instruct Bai et al. (2023).
@@
  • Contrastive Learning: A variant of the SFT model that incorporates disentaglement loss to better distinguish between different feature levels. Detailed training objectives and implementation details are provided in Appendix G.
@@ 5.1.2 Results and Analysis
The experimental results, summarized in Table 1 and 9, reveal several key insights regarding the controllability of seeker simulators. Linguistic features such as verbosity and dialogue length are relatively easy to learn, with all training-based methods substantially outperforming zero-shot baselines. In contrast, psychological features like resistance and self-disclosure are much harder: zero-shot models score around 0.2 in Macro F1, standard training reaches only around 0.3, and only our MoE-based model exceeds 0.4. Feature-level accuracy and correlation scores are in Table 9 and 9.
@@
This validates that our proposed MoE framework autonomously induces the emergent specialization of intrinsic parameter subspaces, effectively partitioning behavioral counseling regimes. Detailed figures are provided in Appendix I.
@@ 5.2.1 Experimental Setup
For each trial, experts were presented with two dialogues generated by different seeker simulators starting from the same initial problem summary. Experts select the dialogue whose seeker utterances better satisfy each evaluation criterion, or choose "Tie" if the quality is indistinguishable. Experts evaluate seeker fidelity along three dimensions: (1) Linguistic Naturalness, (2) Role Authenticity, and (3) Psychological Plausibility, which respectively assess language fluency, role consistency, and emotional coherence. Detailed definitions and instructions are provided in Appendix H and Table 10
@@
We compare our model against three representative seeker simulators. To ensure that each baseline operates at its intended capacity, we adopt the original profile structures and prompting schemes specified in their respective papers. Specifically, we incorporate Eeyore Liu et al. (2025) and ESC-Role Zhao et al. (2024), both of which are fine-tuned on emotional support datasets, and seeker model in ESC-Judge Madani and Srihari (2025), a zero-shot baseline utilizing GPT-4o without additional training.
@@ 5.2.2 Results and Analysis
The results of the expert evaluation are summarized in Table 11. Overall, our simulator consistently outperformed all baseline models across all three dimensions, achieving a win rate of over 65% in every pairwise comparison.
Specifically, our model showed a distinct advantage in linguistic naturalness over Eeyore and ESC-Role, capturing more raw communicative patterns found in real-world emotional support interactions. Furthermore, compared to ESC-Judge, our simulator demonstrated higher role authenticity, as individuals often exhibit uncertainty about their own problems rather than producing overly structured or verbose explanations. Most significantly, the superior performance in psychological plausibility validates that our simulator captures authentic patterns of emotional change, maintaining a coherent emotional trajectory across the dialogue.

### v2

# Stress-Testing Emotional Support Models: Moving from Homogeneous to Diverse Help Seekers

@@ Abstract
Stress-Testing Emotional Support Models:
Moving from Homogeneous to Diverse Help Seekers
Chaewon Heo  Cheyon Jin  Yohan Jothanks: Corresponding author Graduate School of Data Science, Seoul National University {heorshey99, cheyonjin, yohan.jo}@snu.ac.kr
@@ 1 Introduction
Currently, the most prevalent evaluation strategy uses help-seeker simulators to interact with supporter models, and the resulting dialogues are assessed based on metrics such as empathy and fluency Zhao et al. (2024). However, existing seeker simulators compromise evaluation validity due to two critical limitations. First, they fail to reflect the diversity of real-world seekers, representing only a narrow subset of seeker behaviors. Existing simulators primarily portray overly cooperative, "easy" seekers, failing to capture the vast spectrum of real-world behaviors such as advice resistance Yaman (2021) or limited disclosure. Second, existing simulators lack fine-grained controllability. A robust evaluation framework must allow researchers to target specific seeker populations, as this enables the identification of which aspects of a supporter model fail for particular seeker groups, providing actionable signals for model improvement.
To address these shortcomings, we present a highly controllable seeker simulator designed to take specific seeker profiles and simulate their corresponding behaviors. As a foundational component, we define the seeker profile as a combination of nine psychological and linguistic features such as resistance level and verbosity level (Step 1 in Figure 1). To faithfully reflect a broad array of real-world behaviors, we leverage data from Reddit online support groups as the primary training source for our simulator.
To effectively control a seeker profile defined by nine features, we employ a Mixture-of-Experts (MoE) architecture. This architecture dynamically assigns weights to specialized adapters based on the input seeker profile, accurately reflecting distinct behavioral patterns across seeker types.
Validation results highlight the effectiveness of our simulator across three key dimensions: (1) superior profile adherence, successfully maintaining consistent seeker traits throughout the interaction; (2) high behavioral diversity, capturing a wider array of help-seeking patterns compared to existing baselines; and (3) higher fidelity, as confirmed by expert evaluations indicating that the generated responses more closely resemble real-world emotional support interactions.
Using our seeker simulator, we evaluate supporter models through interactive multi-turn dialogues, followed by automatic assessment using emotional support and general conversational skill metrics (Steps 2, 3 in Figure 1). Results indicate that supporter models exhibit varying levels of performance degradation across various seeker populations. Our framework allows researchers to flexibly define seeker populations by selecting appropriate combinations of profile features, enabling targeted and diagnostic evaluation under diverse scenarios.
@@ 2.1 Evaluation of Emotional Support Models
The evaluation of emotional support models has evolved from surface-level fluency to interactional supportive efficacy. Early research evaluated whether generated empathetic responses exhibited emotionally appropriate expressions Santhanam and Shaikh (2019), using lexical metrics (e.g., BLEU, Perplexity) against reference dialogues such as EmpatheticDialogues Rashkin et al. (2019). Since these metrics often fail to capture supportive effectiveness, later work adopted ESConv Liu et al. (2021) as evaluation scenarios and performed human evaluation, where evaluators role-play seekers and rate support quality using predefined criteria (e.g., Identification, Empathy) Zhao et al. (2023); Cheng et al. (2022). However, these human role-play evaluations are constrained by predefined scenarios, fail to reflect real-world interaction dynamics, and are costly to scale. These limitations have driven recent evaluation toward simulator-based evaluation frameworks, where seeker simulators interact with supporter models and automated evaluation is performed on the generated dialogues.
@@ 2.2 Seeker Simulations in Emotional Support
Prior studies have adopted diverse approaches to implement seeker simulators. Prompt-based simulators are given input including specific characteristics, such as Big Five personality traits, or situational mental health contexts Madani and Srihari (2025); Qiu and Lan (2024). Beyond simple prompting, some work employs explicit state tracking or memory modules to regulate the seeker’s internal states Yang et al. (2025); Wang et al. (2025). Another work trains simulators on emotional support datasets using profile cards with demographic attributes, enabling it to learn seeker behaviors associated with each profile Zhao et al. (2024).
Despite conditioning on different seeker prompts, existing simulators inherit the base language model’s verbose and cooperative interaction bias. Moreover, even learned simulators tend to reflect only surface-level profile attributes, failing to induce deeper psychological traits across seeker types. Our simulator addresses these limitations.
@@ 3.1 Data Collection and Preprocessing
We curated a dataset of real-world emotional support interactions from Reddit, specifically targeting subreddits with over 500,000 members, such as r/offmychest and r/mentalhealth. Reddit serves as an ideal source for training a feature-controllable simulator, as its in-the-wild nature encompasses a broad spectrum of informal and raw human expressions often absent in curated datasets.
Unlike traditional approaches that rely on heavily filtered datasets, we intentionally retained informal and raw elements. We included a wide variance in utterance lengths and aggressiveness, addressing the limitations of prior simulators that portray only articulate and cooperative seekers.
@@ 3.2 Feature Taxonomy for Seeker Profiles
To represent the multifaceted characteristics of seekers’ conversational behavior, we define a comprehensive taxonomy consisting of nine distinct features. These are categorized into psychological and linguistic attributes. Detailed category definitions for each feature are provided in Appendix B.
@@ 3.2.1 Psychological Features
We incorporate six psychological features to model seeker behavior. The main coping strategy categorizes seekers into four types, grounded in the multidimensional coping framework Endler and Parker (1994): problem_focused, emotion_focused, avoidant, and maladaptive_behavior. The utterance style captures stylistic variation in seeker expression using three categories (plain, upset, verbose), adapted from the PATIENT-$\Psi$ framework Wang et al. (2024b). To model interactional dynamics, we define a three-level ordinal resistance level based on psychological reactance theory Brehm and Brehm (1981), and a three-level engagement level grounded in prior work on client participation in therapeutic settings Holdsworth et al. (2014). Finally, we include a four-level self-disclosure level derived from Social Penetration Theory Altman and Taylor (1973), as well as seeker reaction proportions following prior work on modeling client reactions in emotional support dialogues Li et al. (2023).
@@ 3.2.2 Linguistic Features
We further define three linguistic attributes that capture diverse structural patterns. The verbosity level is defined as a five-level feature based on the distribution of seeker utterance lengths. We additionally include a binary user profanity flag, detected using the profanity_check library. Finally, the total dialogue turns level captures the overall length of the interaction.
@@ 3.3 Feature Annotation and Validation
To apply the feature taxonomy to our corpus, we employed a hybrid annotation framework combining LLM-based tagging and rule-based extraction.
@@ 3.3.1 Psychological Feature Annotation
To verify the reliability of the LLM-based feature tagging, we conducted a validation study on a sample of 60 dialogues. Inter-annotator agreement (percent agreement) averages 0.57 across psychological features, while human–LLM alignment achieves 0.84 accuracy, indicating reliable alignment. Detailed validation procedures and feature-level validation results are provided in Appendix D.
@@ 3.3.2 Linguistic Feature Extraction
Linguistic features describe the measurable characteristics of the dialogue and are extracted using rule-based methods: verbosity level from seeker utterance length, profanity flag via profanity-check, and total dialogue turns level from turn count. Detailed extraction rules are in Appendix E.
@@ 3.4 Seeker Profile Construction
We construct a final seeker profile by combining all annotated features with the seeker’s main problem. Formally, a seeker profile is represented as:
@@
where each $f_{i}$ denotes an annotated seeker feature and $m$ represents the seeker’s main problem, both expressed in natural language. An example profile is in Figure 8. The resulting dataset consists of 11,066 unique seeker profiles, split into 8,868 training, 1,094 validation, and 1,104 test profiles. The seeker’s main problem is a summary of the first seeker utterance (i.e., original post) that describes the seeker’s situation. The summaries are generated by GPT-4o-mini OpenAI (2024), with detailed prompts provided in Table 12.
Figure 2: Overview of the MoE architecture. A frozen SFT backbone is augmented with multiple low-rank expert adapters at each linear layer. A shared routing network produces a dialogue-level routing vector $\boldsymbol{\alpha}$, which controls expert activation consistently across all layers.
@@ 4 Seeker Simulator Training
To simulate diverse seeker behaviors in emotional support conversations, we employ a two-stage training framework: (1) Supervised Fine-Tuning (SFT) to establish a general conversational backbone, and (2) Mixture-of-Experts (MoE) training to enable controllable, attribute-specific generation.
@@ 4.1 Supervised Fine-Tuning
The first stage establishes a linguistic foundation by capturing general conversational patterns of seekers. We use Llama-3-8B-Instruct as the base model and apply LoRA ($r=16$) to all linear layers. The model is trained to predict the next seeker utterance, given a seeker profile and dialogue history as a next-token prediction task. Upon convergence, the LoRA adapters are merged into the base model weights. This merged model is then entirely frozen, serving as a stable backbone for the subsequent expert-based training stage.
@@ 4.2 MoE Training with Behavioral Routing
To address this challenge, we introduce an explicit routing mechanism that decouples feature control from language modeling. Rather than expecting the model to infer seeker features from text, we provide a structured feature vector to control expert selection through a soft MoE formulation. An overview of the proposed framework is in Figure 2.
@@ 5.1.1 Experimental Setup
  • Zero-shot Base Models: GPT-4.1-mini, Llama-3-8B-Instruct Touvron and others (2024), Qwen-2.5-14B-Instruct Bai and others (2023), GPT-5 Singh et al. (2025), and DeepSeek-V3.2 DeepSeek-AI et al. (2025)
@@
  • Contrastive Learning: A variant of the SFT model that incorporates disentanglement loss to better distinguish between different feature levels. Detailed training objectives and implementation details are provided in Appendix F.
@@ 5.1.2 Results and Analysis
The experimental results, summarized in Table 1 and 14, reveal several key insights regarding the controllability of seeker simulators. Linguistic features such as verbosity and dialogue length are relatively easy to learn, with all training-based methods substantially outperforming zero-shot baselines. In contrast, psychological features like resistance and self-disclosure are much harder: zero-shot models score around 0.2 in Macro F1, standard training reaches only around 0.3, and only our MoE-based model exceeds 0.4. Overall, our model achieves the highest profile adherence across all baselines, including standard zero-shot approaches, frontier reasoning models, and other training-based methods. Feature-level accuracy and correlation scores are in Table 14 and 14.
@@
To identify the mechanism behind improved profile adherence, we analyze whether individual experts specialize in distinct behavioral roles.
Specifically, we interpret expert behavior by first assigning each dialogue to a dominant expert, defined as $e^{*}=\arg\max\boldsymbol{\alpha}$, where $\boldsymbol{\alpha}$ is the routing vector. We visualize the routing distributions $\boldsymbol{\alpha}$ using PCA (Figure 5), where each point is colored by its dominant expert. The clear separation of clusters indicates that the routing network partitions the representation space into distinct regions, providing a structural basis for expert-level interpretation.
We then analyze the conditional distributions $P(\text{feature}\mid e^{*})$ over these assignments. These distributions are visualized as row-normalized heatmaps in Figure 6. As an example, Expert 1 shows strong concentration on high engagement and self-disclosure, along with lower resistance in the heatmap (e.g., engagement: 1.0, resistance: 0.76, self-disclosure level 4: 0.53), corresponding to a highly cooperative and open behavioral pattern.
@@
This suggests that our proposed MoE framework induces the emergent specialization of intrinsic parameter subspaces, effectively partitioning behavioral counseling regimes. Detailed figures are provided in Appendix G.
@@ 5.2.1 Experimental Setup
For each trial, experts were presented with two dialogues generated by different seeker simulators starting from the same initial problem summary. Experts select the dialogue whose seeker utterances better satisfy each evaluation criterion, or choose "Tie" if the quality is indistinguishable. Experts evaluate seeker fidelity along three dimensions: (1) Linguistic Naturalness, (2) Role Authenticity, and (3) Psychological Plausibility, which respectively assess language fluency, role consistency, and emotional coherence. Detailed definitions, instructions, annotator information, and inter-annotator agreement are provided in Appendix H and Table 15.
@@
We compare our model against three representative seeker simulators. To ensure that each baseline operates at its intended capacity, we adopt the original profile structures and prompting schemes specified in their respective papers. Specifically, we incorporate the seeker simulator in ESC-Judge Madani and Srihari (2025), a zero-shot baseline utilizing GPT-4o without additional training, and ESC-Role Zhao et al. (2024) and Eeyore Liu et al. (2025), both of which are fine-tuned on emotional support datasets.
@@ 5.2.2 Results and Analysis
The results of the expert evaluation are summarized in Table 2. Overall, our simulator consistently outperformed all baseline models across all three dimensions, achieving an average win rate of 69.5% across all pairwise comparisons and criteria.
Ours vs. Eeyore 68.9 66.7 71.1
Ours vs. ESC-Judge 68.9 72.2 62.2
Ours vs. ESC-Role 80.0 67.8 67.8
Table 2: Win rates (%) of our simulator against baseline simulators across three fidelity criteria. Win rate is computed as the proportion of wins out of total comparisons, including ties.
Specifically, our model showed a distinct advantage in linguistic naturalness over Eeyore and ESC-Role, capturing more raw communicative patterns found in real-world emotional support interactions. Furthermore, compared to ESC-Judge, our simulator demonstrated higher role authenticity, as individuals often exhibit uncertainty about their own problems rather than producing overly structured or verbose explanations. Most significantly, the superior performance in psychological plausibility validates that our simulator captures authentic patterns of emotional change, maintaining a coherent emotional trajectory across the dialogue. Detailed win/loss/tie breakdowns are provided in Table 16.